MAY I PLEASE GET SOME HELP WITH THESE I REALLY NEED HELP
A regular squre pryramid with slant height 9 and base edge 6
(a)lateral area
(b)the total area
A right cylinder with radius 3 and height 7
(a)lateral area
(b) the total area
Two circles have area of 49pi and 36pi. Find the ratio of their radii.

Answer by Flake(45)

You need to understand the terms "lateral Area" and "Total Area". Once understand that, it should be pretty straight forward.
A regular squre pryramid with slant height 9 and base edge 6
a) Lateral Area for regular pryramid = 1/2*(perimeter of the base)*(slant height)
-----slant height is given, you need to find the perimeter for the square base
b) the total area = lateral area + base area
----- you know the lateral area from "a". You should know the area for a square base
------------------------
A right cylinder with radius 3 and height 7
a) Lateral Area = (height) * (circumference of the base)
---------- height is given. The circumference of a cone is a circle. You need to calculate the circumference of a circle whose radius is 3
b) the total area = lateral area + base areas
----- you know the lateral area from "a". You should how to calculate the area of a circle base.
------ remember a cone has 2 circle bases. One on the bottom the other is the top.
---------------
Two circles have area of 49pi and 36pi. Find the ratio of their radii.
Let R#1 be radius#1.
Let R#2 be radius#2.
Radio = R#1/R#2
---Find R#1: Area of circle=(pi)(R^2)
So 49pi=(pi)(R#1^2)
So 49=R#1^2
So R#1 = sqrt of 49
-----Now repeat to find R#2 and find ratio